# Jan Blumenkamp, Ajay Shankar, Matteo Bettini, Joshua Bird, Amanda Prorok

### v1

@@ 5 Conclusion
This article introduces a versatile omni-directional ground robot system that is based on extensive modification and enhancement of the original DJI RoboMaster S1 platform. We significantly extend the capabilities of the base platform by providing multiple options for more flexible and powerful compute solutions, onboard sensors, model-based control, and sim-to-real capabilities for distributed RL policies. After an intensive research and development pipeline over four years, our full platform (hardware and software) now serves as a highly reliable testbed well-suited for a wide range of experiments. We showcase its capabilities through four new evaluations, and refer to five case studies from prior work that point to its proven track record.

### v2

@@ 5 Conclusion
This article introduces a versatile omnidirectional ground robot system that is based on extensive modification and enhancement of the original DJI RoboMaster S1 platform. We significantly extend the capabilities of the base platform by providing multiple options for more flexible and powerful compute solutions, onboard sensors, model-based control, and sim-to-real capabilities for distributed RL policies. After an intensive research and development pipeline over four years, our full platform (hardware and software) now serves as a highly reliable testbed well-suited for a wide range of experiments. We showcase its capabilities through four new evaluations with fleet sizes of up to eight robots, and refer to six case studies from prior work that point to its proven track record.